# FAILURE OF LICHNEROWICZ-TYPE RESULT IN PARABOLIC GEOMETRIES OF REAL RANK AT LEAST 33

### v3

@@ 2.4 Harmonic Curvature
The Killing form induces an isomorphism $\mathfrak{g}_{-}^{*}\cong\mathfrak{g}_{+}=\mathfrak{p}_{+}$. Let $\mathfrak{g}\curvearrowright V$ be an action. Define $C_{k}(\mathfrak{g}_{+},V):=\bigwedge^{k}\mathfrak{g}_{+}\otimes V$ and $C^{k}(\mathfrak{g}_{-},V):=\bigwedge^{k}\mathfrak{g}_{-}^{*}\otimes V.$ There is an isomorphism
$C_{k}(\mathfrak{g}_{+},V)\cong C^{k}(\mathfrak{g}_{-},V)$
of $\mathfrak{g}_{0}$ modules. Let $\partial:C_{*}(\mathfrak{g}_{+},V)\to C_{*}(\mathfrak{g}_{+},V)$ be the boundary map for Lie algebra homology, and let $\partial^{*}:C^{*}(\mathfrak{g}_{+},V)\to C^{*}(\mathfrak{g}_{+},V)$ be the coboundary map for Lie algebra cohomology. Both are $\mathfrak{g}_{0}$-equivariant. Then we may define the algebraic Laplacian operator $\Delta:C^{*}(\mathfrak{g}_{-},V)\to C^{*}(\mathfrak{g}_{-},V)$ by $\Delta:=\partial\partial^{*}+\partial^{*}\partial,$ using the identifications between chains and cochains where appropriate. Elements of $\ker\Delta$ are called harmonic. The operators $\partial$ and $\partial^{*}$ are adjoint relative to a certain positive definite inner product, and the algebraic Hodge theory of these spaces implies $\ker\Delta=\ker\partial\cap\ker\partial^{*}$, and the existence of canonical $\mathfrak{g}_{0}$ equivariant isomorphisms
$H_{*}(\mathfrak{g}_{+},V)\cong\ker\Delta\cong H^{*}(\mathfrak{g}_{-},V).$
@@
$\Omega_{p}\in\bigwedge^{2}(\mathfrak{g}/\mathfrak{p})^{*}\otimes\mathfrak{g}\cong C_{2}(\mathfrak{g}_{+},\mathfrak{g})\cong C^{2}(\mathfrak{g}_{-},\mathfrak{g})$
is contained in positive homogeneity, so $\Omega_{p}\in C_{2}(\mathfrak{g}_{+},V)_{+}\cong C^{2}(\mathfrak{g}_{-},\mathfrak{g})_{+}.$ A parabolic Cartan geometry is called normal if $\Omega_{p}\in\ker\partial^{*}$ for every $p\in\mathcal{G}.$ Given a Klein pair $(G,P),$ it is often natural to consider only the regular, normal Cartan geometries modeled on $(G,P).$ If $\Omega$ is normal, it determines at every point an equivalence class $\overline{\Omega}_{p}\in H_{2}(\mathfrak{g}_{+},\mathfrak{g})\cong H^{2}(\mathfrak{g}_{-},\mathfrak{g}).$ Regularity in combination with normality implies $\overline{\Omega}_{p}\in H^{2}(\mathfrak{g}_{-},\mathfrak{g})_{+}$. Since $\Omega_{p}$ and thus $\overline{\Omega}_{p}$ are $P$-equivariant, there is a section associated to $\overline{\Omega}$ in $\mathcal{G}\times_{P}H^{2}(\mathfrak{g}_{-},\mathfrak{g})$ called the harmonic curvature of $\mathcal{G}.$ Harmonic curvature is a complete invariant obstructing flatness of a Cartan geometry. The following is a consequence of Theorem 3.1.12 of [2].
@@
For a regular, normal parabolic geometry, $\overline{\Omega}=0$ implies global flatness.
@@
2.5 Borel-Weil-Bott
Let $\mathfrak{g}$ be complex semisimple with a fixed parabolic subalgebra. This induces a grading on $\mathfrak{g}$ and a decomposition $\mathfrak{g}=\mathfrak{g}_{-}\oplus\mathfrak{g}_{0}\oplus\mathfrak{g}_{+}.$ For $\alpha_{i}\in\Delta^{0},$ let $s_{i}:\mathfrak{h}^{*}\to\mathfrak{h}^{*}$ be the reflection about the hyperplane orthogonal to $\alpha_{i}$. Let $W_{\mathfrak{p}}$ be the subgroup of the Weyl group $W$ generated by $s_{i}$ for $\I.\alpha_{i}\in\Delta^{0}\backslash$ I. Then define $\W.W^{\mathfrak{p}}=W_{\mathfrak{p}}\backslash$ W. We will identify $W^{\mathfrak{p}}$ with a canonical set of representatives of these right cosets, namely the unique representative in each coset of minimal length. Next, define $W^{\mathfrak{p}}(i)\subset W^{\mathfrak{p}}$ as the subset of coset representatives having length $i.$ Given $\alpha_{i}\in\Delta^{0},$ define $N(i)\subset\Delta^{0}$ to be the subset connected to $\alpha_{i}$ by an edge in the Dynkin diagram.
@@ 2.5 Borel-Weil-Bott
$W^{\mathfrak{p}}(2)=\{(ij):i\in I,j\in N(i)\cup I\}.$
Remark 2.5.
It is straightforward to see that the left side is contained in the right side. If $\I\alpha_{i}\in\Delta^{0}\backslash$ I then $W_{\mathfrak{p}}(ij)=W_{\mathfrak{p}}(j),$ so $(ij)$ is not minimal in its right $W_{\mathfrak{p}}$ coset and $(ij)\not\in W^{\mathfrak{p}}.$ If $\alpha_{j}\in\Delta^{0}\backslash(N(i)\cup I)$ then $W_{\mathfrak{p}}(ij)=W_{\mathfrak{p}}(ji)=W_{\mathfrak{p}}(i)$, so $(ij)$ is not minimal in its right $W_{\mathfrak{p}}$ coset. Then $(ij)\not\in W^{\mathfrak{p}}.$
Let $\lambda_{1},\cdots,\lambda_{n}$ be the fundamental weights of $\mathfrak{g}.$ A weight $\gamma=\sum\gamma^{i}\lambda_{i}$ is called $\mathfrak{g}_{0}$-dominant if $\gamma^{i}\geq 0$ for every $\I.i\in\Delta^{0}\backslash$ I. Equivalently, $\gamma|_{\mathfrak{h}\cap\mathfrak{g}_{0}^{ss}}$ is $\mathfrak{g}_{0}^{ss}$-dominant. If $\gamma$ is a $\mathfrak{g}_{0}$-dominant integral weight, let $V^{\gamma}$ be the irreducible representation of lowest weight $\gamma.$ Let $\rho:=\sum\lambda_{i}$ be the Weyl vector. For $w\in W$ and $\lambda\in\mathfrak{h}^{*},$ define $w\cdot\lambda:=w(\lambda+\rho)-\rho.$ For $w\in W^{\mathfrak{p}}$ and a $\mathfrak{g}$-dominant weight $\lambda$, $w\cdot\lambda$ is always $\mathfrak{g}_{0}$-dominant. Let $\Phi_{w}=w(\Delta^{-})\cap\Delta^{+}.$ Then $|\Phi_{w}|=l(w)$ and $\sum_{\alpha\in\Phi_{w}}\alpha=\rho-w(\rho).$
Theorem 2.6 (Borel-Weil-Bott).
@@
    The harmonic representative of lowest weight $w\cdot\gamma$ in $H_{\mathbb{C}}^{i}(\mathfrak{g}_{-},V^{\gamma})$ is given by
@@
Proposition 2.7.
If $i\neq j$ then $\Phi_{(ij)}=\{\alpha_{i},s_{i}(\alpha_{j})\}.$
Now suppose $\mathfrak{g}$ is simple and consider the adjoint representation $\mathfrak{g}\curvearrowright\mathfrak{g}$ having highest root $\mu$ and lowest root $\mu.$ It follows from part (b) of Borel-Weil-Bott that harmonic representatives of the lowest weight vectors of $H_{\mathbb{C}}^{2}(\mathfrak{g}_{-},\mathfrak{g})$ are of the form
@@
having $\mathfrak{g}_{0}$ weight
$\sum_{\alpha\in\Phi_{w}}\alpha-w(\mu)=\rho-w(\rho)-w(\mu)=-w\cdot\mu.$
We define
@@
These are the Weyl group elements corresponding to $\mathfrak{g}_{0}$ irreducible representations with vectors of positive homogeneity in the harmonic curvature module $H_{\mathbb{C}}^{2}(\mathfrak{g}_{-},\mathfrak{g})$, so
@@
If $\mathfrak{g}\curvearrowright V$ is a real Lie algebra representation, we call it complex if $V$ admits a complex structure compatible with the action, that is an endomorphism $J:V\to V$ with $J^{2}=-1$ so that $X\cdot(JV)=J(X\cdot V)$ for each $X\in\mathfrak{g}.$ Otherwise, we call it noncomplex. Similarly, we call a real Lie algebra complex or noncomplex when its adjoint representation is complex or noncomplex, respectively.
@@
    If $\mathfrak{g}\curvearrowright V$ is noncomplex then
    $H^{*}_{\mathbb{C}}(\mathfrak{g}_{-}^{\mathbb{C}},V^{\mathbb{C}})\cong H_{\mathbb{R}}^{*}(\mathfrak{g}_{-},V)^{\mathbb{C}}.$
    Furthermore, this isomorphism is the restriction of the isomorphism $C^{*}_{\mathbb{C}}(\mathfrak{g}^{\mathbb{C}},V^{\mathbb{C}})\cong C^{*}_{\mathbb{R}}(\mathfrak{g},V)^{\mathbb{C}}$ to harmonic elements.
@@
    If $\mathfrak{g}\curvearrowright V^{\mathbb{R}}$ is complex then
    $H^{*}_{\mathbb{C}}(\mathfrak{g}_{-}^{\mathbb{C}},V)\cong H^{*}_{\mathbb{R}}(\mathfrak{g}_{-},V^{\mathbb{R}}).$
    Furthermore, this isomorphism is the restriction of the isomorphism $C^{*}_{\mathbb{C}}(\mathfrak{g}_{-}^{\mathbb{C}},V)\cong C^{*}_{\mathbb{R}}(\mathfrak{g}_{-},V^{\mathbb{R}})$ to harmonic elements.
Given representations $\mathfrak{g}\curvearrowright V,$ $\mathfrak{g}^{\prime}\curvearrowright V^{\prime}$, take the representation $\mathfrak{g}\oplus\mathfrak{g}^{\prime}\curvearrowright V\otimes V^{\prime}$ given by
$(X,X^{\prime})\cdot(v\otimes v^{\prime})=(X\cdot v)\otimes v^{\prime}+v\otimes(X^{\prime}\cdot v^{\prime}),$
and denote it $V\boxtimes V^{\prime}.$ The 0-graded part of $\mathfrak{g}\oplus\mathfrak{g}^{\prime}$ is $\mathfrak{g}_{0}\oplus\mathfrak{g}^{\prime}_{0}.$ There is a $\mathfrak{g}_{0}\oplus\mathfrak{g}_{0}^{\prime}$-module isomorphism
$\begin{split}C^{*}(\mathfrak{g}_{-},V)\otimes C^{*}(\mathfrak{g}^{\prime}_{-},V^{\prime})&\cong(\Lambda^{*}(\mathfrak{g}_{-})\otimes V)\otimes(\Lambda^{*}(\mathfrak{g}^{\prime}_{-})\otimes V^{\prime})\\ &\cong(\Lambda^{*}(\mathfrak{g}_{-})\otimes\Lambda^{*}(\mathfrak{g}^{\prime}_{-}))\otimes(V\otimes V^{\prime})\\ &\cong\Lambda^{*}(\mathfrak{g}_{-}\oplus\mathfrak{g}^{\prime}_{-})\otimes(V\otimes V^{\prime})\\ &\cong C^{*}(\mathfrak{g}_{-}\oplus\mathfrak{g}^{\prime}_{-},V\otimes V^{\prime}).\end{split}$ (1)
acting by
$(\eta_{1}\wedge\dots\eta_{k}\otimes v)\otimes(\eta^{\prime}_{1}\wedge\dots\wedge\eta^{\prime}_{n}\otimes v^{\prime})\mapsto\eta_{1}\wedge\dots\wedge\eta_{k}\wedge\eta^{\prime}_{1}\wedge\dots\wedge\eta_{l}\otimes(v\otimes v^{\prime}).$ (2)
Using the Borel-Weil-Bott theorem, it is possible to prove the following result.
@@
If $\mathfrak{g},\mathfrak{g}^{\prime}$ are complex semisimple with representations $\mathfrak{g}\curvearrowright V$ and $\mathfrak{g}^{\prime}\curvearrowright V^{\prime}$ then
$H_{\mathbb{C}}^{*}(\mathfrak{g},V)\boxtimes H_{\mathbb{C}}^{*}(\mathfrak{g}^{\prime},V^{\prime})\cong H^{*}(\mathfrak{g}\oplus\mathfrak{g}^{\prime},V\boxtimes V^{\prime})$ (3)
as $\mathfrak{g}_{0}\oplus\mathfrak{g}^{\prime}_{0}$-modules, and the map is given on harmonic representatives by equation (2).
Corollary 2.10.
Let $\mathfrak{p}\leq\mathfrak{g}$ be a complex parabolic subalgebra. Then there is a $\mathfrak{g}_{0}^{\mathbb{R}}$-equivariant injection $H^{2}_{\mathbb{C}}(\mathfrak{g}_{-},\mathfrak{g})\hookrightarrow H^{2}_{\mathbb{R}}(\mathfrak{g}_{-}^{\mathbb{R}},\mathfrak{g}^{\mathbb{R}}).$ Furthermore, this injection is the restriction of the map $C^{2}_{\mathbb{C}}(\mathfrak{g}_{-},\mathfrak{g})\hookrightarrow C^{2}_{\mathbb{R}}(\mathfrak{g}_{-}^{\mathbb{R}},\mathfrak{g}^{\mathbb{R}})$ to harmonic elements.
@@
Since $\mathfrak{g}_{-}$ acts on $\mathfrak{g}$ by complex endomorphisms, the action $\mathfrak{g}_{-}^{\mathbb{R}}\curvearrowright\mathfrak{g}^{\mathbb{R}}$ is complex. Since $\mathfrak{g}_{-}$ is a complex Lie algebra, $(\mathfrak{g}_{-}^{\mathbb{R}})^{\mathbb{C}}\cong\mathfrak{g}_{-}\oplus\mathfrak{g}_{-}.$ In the natural extension $(\mathfrak{g}_{-}^{\mathbb{R}})^{\mathbb{C}}\curvearrowright\mathfrak{g}$ of the action, the first $\mathfrak{g}_{-}$ factor acts by the adjoint action, while the second $\mathfrak{g}_{-}$ factor acts trivially. Then by Proposition 2.8(b),
$H_{\mathbb{R}}^{*}(\mathfrak{g}_{-}^{\mathbb{R}},\mathfrak{g}^{\mathbb{R}})$ $\cong H^{*}_{\mathbb{C}}((\mathfrak{g}_{-}^{\mathbb{R}})^{\mathbb{C}},\mathfrak{g})$
$\cong H^{*}_{\mathbb{C}}(\mathfrak{g}_{-}\oplus\mathfrak{g}_{-},\mathfrak{g})$
$\cong H^{*}_{\mathbb{C}}(\mathfrak{g}_{-},\mathfrak{g})\boxtimes H^{*}_{\mathbb{C}}(\mathfrak{g}_{-},\mathbb{C}),$
so
$\begin{split}H^{2}_{\mathbb{R}}(\mathfrak{g}_{-}^{\mathbb{R}},\mathfrak{g}^{\mathbb{R}})&\cong H^{2}_{\mathbb{C}}(\mathfrak{g}_{-},\mathfrak{g})\boxtimes H^{0}_{\mathbb{C}}(\mathfrak{g}_{-},\mathbb{C})\\ &\oplus H^{1}_{\mathbb{C}}(\mathfrak{g}_{-},\mathfrak{g})\boxtimes H^{1}_{\mathbb{C}}(\mathfrak{g}_{-},\mathbb{C})\\ &\oplus H^{0}_{\mathbb{C}}(\mathfrak{g}^{-},\mathfrak{g})\boxtimes H^{2}_{\mathbb{C}}(\mathfrak{g}_{-},\mathbb{C}).\end{split}$ (4)
Since $H^{0}_{\mathbb{C}}(\mathfrak{g}_{-},\mathbb{C})\cong\mathbb{C}$, it follows that
$H^{2}_{\mathbb{C}}(\mathfrak{g}_{-},\mathfrak{g})\boxtimes H^{0}_{\mathbb{C}}(\mathfrak{g}_{-},\mathbb{C})\cong H^{2}_{\mathbb{C}}(\mathfrak{g}_{-},\mathfrak{g}).$

### v4

@@ 3.1 Harmonic Curvature
The Killing form induces an isomorphism $\mathfrak{g}_{-}^{*}\cong\mathfrak{p}_{+}$. Let $\mathfrak{g}\curvearrowright V$ be an action. Define $C_{k}(\mathfrak{p}_{+},V):=\bigwedge^{k}\mathfrak{p}_{+}\otimes V$ and $C^{k}(\mathfrak{g}_{-},V):=\bigwedge^{k}\mathfrak{g}_{-}^{*}\otimes V.$ There is an isomorphism
$C_{k}(\mathfrak{p}_{+},V)\cong C^{k}(\mathfrak{g}_{-},V)$
of $\mathfrak{g}_{0}$ modules. Let $\partial:C_{*}(\mathfrak{p}_{+},V)\to C_{*}(\mathfrak{p}_{+},V)$ be the boundary map for Lie algebra homology, and let $\partial^{*}:C^{*}(\mathfrak{p}_{+},V)\to C^{*}(\mathfrak{p}_{+},V)$ be the coboundary map for Lie algebra cohomology. Both are $\mathfrak{g}_{0}$-equivariant. Then we may define the algebraic Laplacian operator $\Delta:C^{*}(\mathfrak{g}_{-},V)\to C^{*}(\mathfrak{g}_{-},V)$ by $\Delta:=\partial\partial^{*}+\partial^{*}\partial,$ using the identifications between chains and cochains where appropriate. Elements of $\ker\Delta$ are called harmonic. The operators $\partial$ and $\partial^{*}$ are adjoint relative to a certain positive definite inner product, and the algebraic Hodge theory of these spaces implies $\ker\Delta=\ker\partial\cap\ker\partial^{*}$, and the existence of canonical $\mathfrak{g}_{0}$-equivariant isomorphisms
$H_{*}(\mathfrak{p}_{+},V)\cong\ker\Delta\cong H^{*}(\mathfrak{g}_{-},V).$
@@
$\Omega_{p}\in\bigwedge^{2}(\mathfrak{g}/\mathfrak{p})^{*}\otimes\mathfrak{g}\cong C_{2}(\mathfrak{p}_{+},\mathfrak{g})\cong C^{2}(\mathfrak{g}_{-},\mathfrak{g})$
is contained in positive homogeneity, so $\Omega_{p}\in C_{2}(\mathfrak{p}_{+},\mathfrak{g})_{+}\cong C^{2}(\mathfrak{g}_{-},\mathfrak{g})_{+}.$ A parabolic Cartan geometry is called normal if $\Omega_{p}\in\ker\partial^{*}$ for every $p\in\mathcal{G}.$ Given a Klein pair $(G,P),$ it is often natural to consider only the regular, normal Cartan geometries modeled on $(G,P).$ If $\Omega$ is normal, it determines at every point an equivalence class $\overline{\Omega}_{p}\in H_{2}(\mathfrak{p}_{+},\mathfrak{g})\cong H^{2}(\mathfrak{g}_{-},\mathfrak{g}).$ Regularity in combination with normality implies $\overline{\Omega}_{p}\in H^{2}(\mathfrak{g}_{-},\mathfrak{g})_{+}$. The function $\overline{\Omega}$ is called the harmonic curvature of the Cartan geometry. Harmonic curvature is a complete invariant obstructing flatness of a Cartan geometry. The following is a consequence of Theorem 3.1.12 of [2].
@@
For a regular, normal parabolic geometry, $\overline{\Omega}=0$ implies flatness of the Cartan geometry.
@@
3.2 Kostant’s Borel-Weil-Bott
Let $\mathfrak{g}$ be complex semisimple with a fixed parabolic subalgebra. This induces a grading on $\mathfrak{g}$ and an associated decomposition $\mathfrak{g}=\mathfrak{g}_{-}\oplus\mathfrak{g}_{0}\oplus\mathfrak{p}_{+}$ into the negative, zero, and positive graded components. For $\alpha_{i}\in\Delta^{0},$ let $s_{i}:\mathfrak{h}^{*}\to\mathfrak{h}^{*}$ be the reflection about the hyperplane orthogonal to $\alpha_{i}$. Let $W^{\mathfrak{p}}\subset W$ be the Hasse diagram associated to $\mathfrak{p}$, as explained in [2], and let $W^{\mathfrak{p}}(i)\subset W^{\mathfrak{p}}$ be the subset of elements of length $i.$ Given $\alpha_{i}\in\Delta^{0},$ define $N(i)\subset\Delta^{0}$ to be the subset connected to $\alpha_{i}$ by an edge in the Dynkin diagram. The following is Recipe 3 of [9].
@@ 3.2 Kostant’s Borel-Weil-Bott
The Hasse diagram elements of length $2$ are characterized by
$W^{\mathfrak{p}}(2)=\{(ij):\alpha_{i}\in I,\alpha_{j}\in N(i)\cup I\}.$
Let $\lambda_{1},\cdots,\lambda_{n}$ be the fundamental weights of $\mathfrak{g}.$ A weight $\gamma=\sum\gamma^{i}\lambda_{i}$ is called $\mathfrak{g}_{0}$-dominant if $\gamma^{i}\geq 0$ for every $\I.i\in\Delta^{0}\backslash$ I. Equivalently, $\gamma|_{\mathfrak{h}\cap\mathfrak{g}_{0}^{ss}}$ is $\mathfrak{g}_{0}^{ss}$-dominant. If $\gamma$ is a $\mathfrak{g}_{0}$-dominant integral weight, let $V^{\gamma}$ be the irreducible representation of lowest weight $\gamma.$ Let $\rho:=\sum\lambda_{i}$ be the Weyl vector. For $w\in W$ and $\lambda\in\mathfrak{h}^{*},$ define $w\cdot\lambda:=w(\lambda+\rho)-\rho.$ For $w\in W^{\mathfrak{p}}$ and a $\mathfrak{g}$-dominant weight $\lambda$, $w\cdot\lambda$ is always $\mathfrak{g}_{0}$-dominant. Let $\Phi_{w}=w(\Delta^{-})\cap\Delta^{+}.$ In particular, $\Phi_{(ij)}=\{\alpha_{i},s_{i}(\alpha_{j})\}.$ The following is a consequence of Theorem 3.3.5 in [2].
Theorem 3.3 (Kostant’s Borel-Weil-Bott).
@@
    A harmonic representative of lowest weight $w\cdot\gamma$ in $H_{\mathbb{C}}^{i}(\mathfrak{g}_{-},V^{\gamma})$ is given by
@@
Now suppose $\mathfrak{g}$ is simple with highest root $\mu$. Then the adjoint representation $\mathfrak{g}\curvearrowright\mathfrak{g}$ has lowest weight $\mu.$ It follows from part (b) of Kostant’s Borel-Weil-Bott that harmonic representatives of the lowest weight vectors of $H_{\mathbb{C}}^{2}(\mathfrak{g}_{-},\mathfrak{g})$ are of the form
@@
of $\mathfrak{g}_{0}$ weight $w\cdot\mu.$ We define
@@
These are the Weyl group elements corresponding to $\mathfrak{g}_{0}$ irreducible subrepresentations in $H_{\mathbb{C}}^{2}(\mathfrak{g}_{-},\mathfrak{g})$ of positive homogeneity. Therefore
@@
A morphism of representations $(\mathfrak{g},V)\to(\mathfrak{g}^{\prime},V^{\prime})$ is a Lie algebra homomorphism $\phi:\mathfrak{g}\to\mathfrak{g}^{\prime}$ and a linear map $\psi:V\to V^{\prime}$ such that $\psi(X\cdot v)=\phi(X)\cdot\psi(v).$ A morphism of representations induces maps $C^{*}(\mathfrak{g},V)\to C^{*}(\mathfrak{g}^{\prime},V^{\prime})$ and $C_{*}(\mathfrak{g},V)\to C_{*}(\mathfrak{g}^{\prime},V^{\prime})$ which are $\partial$ and $\partial^{*}$-equivariant, respectively. In particular this holds if $\mathfrak{g}^{\prime}=\mathfrak{g}$ with $\phi=\mathrm{Id}_{\mathfrak{g}}$ and $\psi:V\to V^{\prime}$ is $\mathfrak{g}$-equivariant. The following is a slight extension of parts (1) and (2) of Proposition 3.3.6 in [2].
@@
    If $\mathfrak{g}$ acts on a complex vector space $V$ by complex endomorphisms,
    $H^{*}_{\mathbb{R}}(\mathfrak{g},V)\cong H^{*}_{\mathbb{C}}(\mathfrak{g}^{\mathbb{C}},V).$
    Furthermore, map is the restriction of the isomorphism $C^{*}_{\mathbb{R}}(\mathfrak{g},V)\cong C^{*}_{\mathbb{C}}(\mathfrak{g}^{\mathbb{C}},V)$ to harmonic elements.
@@
    If $\mathfrak{g}$ acts on a real vector space $V$ by real endomorphisms then
    $H_{\mathbb{R}}^{*}(\mathfrak{g},V)^{\mathbb{C}}\cong H^{*}_{\mathbb{C}}(\mathfrak{g}^{\mathbb{C}},V^{\mathbb{C}}).$
    Furthermore, this isomorphism is the restriction of the isomorphism $C^{*}_{\mathbb{R}}(\mathfrak{g},V)^{\mathbb{C}}\cong C^{*}_{\mathbb{C}}(\mathfrak{g}^{\mathbb{C}},V^{\mathbb{C}})$ to harmonic elements.
Proof.
  1. (a)
    In the following diagram, the top isomorphism replaces real tensor products with complex ones, while the bottom isomorphism passes to the unique $\mathbb{C}$-linear extension. The left and right isomorphism are induced by the real and complex Killing forms on $\mathfrak{g}$ and $\mathfrak{g}^{\mathbb{C}}$ respectively. The reader may check that the diagram is commutative.
[…]
    The top isomorphism on chains is $\partial^{*}$-equivariant, while the bottom isomorphism on cochains is $\partial$-equivariant. Identifying chains and cochains, it follows from the definition $\Delta=\partial\partial^{*}+\partial^{*}\partial$ that the two isomorphisms are $\Delta$-equivariant. It follows that harmonic cochains in $C_{\mathbb{R}}^{*}(\mathfrak{g},V)$ are mapped bijectively onto harmonic cochains in $C_{\mathbb{C}}^{*}(\mathfrak{g}^{\mathbb{C}},V)$.
  2. (b)
    The following diagram commutes. The top and bottom arrows are $\partial^{*}$ and $\partial$-equivariant because $V\to V^{\mathbb{C}}$ is $\mathfrak{g}$-equivariant.
    $C_{*}^{\mathbb{R}}(\mathfrak{g},V)$$C_{*}^{\mathbb{R}}(\mathfrak{g},V^{\mathbb{C}})$$C^{*}_{\mathbb{R}}(\mathfrak{g},V)$$C^{*}_{\mathbb{R}}(\mathfrak{g},V^{\mathbb{C}})$$\scriptstyle{\cong}\scriptstyle{\cong}$
    Therefore the following diagram is also commutative with top and bottom arrows $\partial^{*}$ and $\partial$-equivariant.
[…]
    It follows that harmonic cochains in $C^{*}_{\mathbb{R}}(\mathfrak{g},V)^{\mathbb{C}}$ are taken to harmonic cochains in $C^{*}_{\mathbb{R}}(\mathfrak{g},V^{\mathbb{C}}).$ Composing with the canonical isomorphism to $C^{*}_{\mathbb{C}}(\mathfrak{g}^{\mathbb{C}},V^{\mathbb{C}})$ takes harmonic cochains bijectively onto harmonic cochains by part (a).
@@
The forgetful map $C^{*}_{\mathbb{C}}(\mathfrak{g}_{-},V)\to C^{*}_{\mathbb{R}}(\mathfrak{g}_{-},\mathfrak{g})$ takes harmonic elements to harmonic elements. Consequently, there is a $\mathfrak{g}_{0}$-equivariant injection $H^{*}_{\mathbb{C}}(\mathfrak{g}_{-},V)\to H^{*}_{\mathbb{R}}(\mathfrak{g}_{-},V).$
@@
The following diagram is commutative.
[…]
As a result, the first two squares of the following diagram are commutative. We checked the commutativity of the third square in the proof of Proposition 3.4 (a).
[…]
The composite of the bottom arrows is the forgetful map. Let $\overline{\mathfrak{p}}_{+}$ and $\overline{\mathfrak{g}_{-}}$ act trivially on $V.$ The first two top arrows are $\partial^{*}$-equivariant because $(\mathfrak{p}_{+},V)\to(\mathfrak{p}_{+}\oplus\overline{\mathfrak{p}_{+}},V)$ and $(\mathfrak{p}_{+}\oplus\overline{\mathfrak{p}_{+}},V)\to(\mathfrak{p}_{+}^{\mathbb{C}},V)$ are morphisms of representations. We checked that the third map is $\partial^{*}$-equivariant in the proof of Proposition 3.4 (a) The bottom arrows are $\partial$-equivariant for similar reasons. The claim follows. ∎